If the mappings f and g are given by f = {(1, 2), (3, 5), (4, 1)} and g = {(2, 3), (5, 1), (1, 3)}, write f o g. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,
f = {(1, 2), (3, 5), (4, 1)} and g = {(2, 3), (5, 1), (1, 3)} Now,
fog (2) = f(g(2)) = f(3) = 5
fog (5) = f(g(5)) = f(1) = 2
fog (1) = f(g(1)) = f(3) = 5
Thus,
fog = {(2, 5), (5, 2), (1, 5)}

Thanks